# New determination of the neutrino hadronic production cross sections from GeV to beyond PeV energies

### v1

@@ Abstract
The flux of astrophysical neutrinos is now measured with unprecedented accuracy and on several decades of energy spectrum. Their origin traces back to hadronic collisions between protons and nuclei in the cosmic rays with hydrogen and helium in the target gas. To accurately interpret the data, a precise determination of the underlying cross sections is therefore mandatory. We present a new evaluation of the neutrino production cross section from $p+p$ collisions, building on our previous analysis of the production cross section for $\pi^{\pm}$, $K^{\pm}$, and minor baryonic and mesonic channels. Cross sections for scatterings involving nuclei heavier than protons are also derived. The novelty of our approach is the analytical description of the Lorentz invariant cross section $\sigma_{\rm inv}$, and the fit of the model to the available accelerator data. We work with neutrino energies from $10$ GeV to $10^{7}$ GeV, and, correspondingly, to incident proton (nuclei) energies from $10$ GeV to $10^{9}$ GeV (GeV/n). We obtain the total differential cross section, $d\sigma(p+p\rightarrow\nu+X)/dE_{\nu}$ as a function of neutrino and proton energies, with an estimated uncertainty of 5% for neutrino energies below 100 GeV, increasing to 10% above TeV energies. Predictions are given for $\nu_{e},\nu_{\mu},\bar{\nu_{e}}$ and $\bar{\nu_{\mu}}$. A comparison with state-of-the-art cross sections, all relying on Monte Carlo generators, is also presented. To facilitate the use by the community, we provide numerical tables and a script for accessing our energy-differential cross sections.
@@ I Introduction
In this paper, we present a new accurate model that relies primarily on an analytic prescription fitted to the available experimental data, with the aim of determining their correct dependence on kinematic variables and of robustly quantifying the modeling uncertainties. The dominant production channel for $\nu_{e}$, $\nu_{\mu}$, $\bar{\nu_{e}}$ and $\bar{\nu_{\mu}}$ is the decay of $\pi^{\pm}$ mesons, for which we build upon the analysis performed in ODDK22. In addition, we carefully model the production cross sections of $K$ mesons and $\Lambda$ baryons, which contribute as well to the $\nu$ flux through their decay into $\pi^{\pm}$. Our strategy closely follows the approach adopted in ODDK22 for deriving cross sections relevant to the secondary production of CR electrons and positrons.
@@ III $\nu$ from $p+p\rightarrow\pi^{\pm}+X$ collisions
The $\pi^{\pm}$ channel for $\nu$ production corresponds to the first two rows in Fig. 1(left) for $\nu_{\mu}$ and to the first row in Fig. 1(right) for $\nu_{e}$. Given its relevance, it is crucial to obtain precise data covering a wide kinematic phase space for the reaction $p+p\rightarrow\pi^{\pm}+X$. As discussed in ODDK22, measurements of $\sigma_{\rm inv}(p+p\rightarrow\pi^{\pm}+X)$ have been collected by various accelerator and collider experiments, spanning large portions of the kinematic phase space. In particular, NA49 Alt et al. (2005) and NA61 Aduszkiewicz et al. (2017) provide data at low $\sqrt{s}$, while ALICE Aamodt (2011) and CMS Chatrchyan et al. (2012); Sirunyan et al. (2017) cover the high-$\sqrt{s}$ regime. In this work we do not perform a new fit to the particle production data (since no new data have appeared); instead, we use the results for $\sigma_{\rm inv}$ obtained in ODDK22 to start the calculation of $\nu$ yields. The expressions for $\sigma_{\rm inv}$ are given in Eqs. (7)–(10) and the parameters for $\pi^{\pm}$ are listed in Table II of ODDK22. We remark that the fit performed in ODDK22 is based on an analytical formula for $\sigma_{\rm inv}$, directly calibrated on the data and their uncertainties. The highest projectile energy on which we tested our model is $T_{p}=9\times 10^{7}$ GeV, corresponding to $\sqrt{s}=13$ TeV pion data from CMS Sirunyan et al. (2017). Beyond this limit, our model can only be extrapolated. However, care must be taken when treating purely statistical uncertainties at high $\sqrt{s}$. As explained in ODDK22, our model of $\sigma_{\rm inv}$ relies on a fit on NA49 and NA61 $\pi^{\pm}$ data at $\sqrt{s}=17.3$ GeV, which provide broad coverage in both $p_{T}$ and $x_{R}$, thereby constraining in a solid way the dependence of the model on these variables. The scaling of the cross section at higher energies was instead determined by fits to CMS and ALICE data at $\sqrt{s}>900$ GeV (corresponding to $T_{p}=4.3\times 10^{5}$ GeV in the LAB frame). These measurements, however, were performed at mid-rapidity and cover only a limited range in $p_{T}$, with no data available at forward rapidities. Consequently, our model implies an extrapolation of the cross section in the forward-rapidity region at high $\sqrt{s}$. To quantify the impact of this incomplete phase-space coverage, we will compare our results with predictions from other state-of-the-art Monte Carlo models, thereby estimating the overall uncertainties across different cross-section parametrizations.
@@
In Fig. 2, we show the differential cross section for the production of $\nu_{\mu}$, $\bar{\nu}_{\mu}$, $\nu_{e}$, and $\bar{\nu}_{e}$ from $\pi^{\pm}$ decays in $p+p$ collisions. Results are provided for different incident kinetic proton energies $T_{p}$ as a function of $E_{\nu}$, for all $\nu$ flavors. The uncertainties range from $5\%$ to $20\%$ over most of the energy range, except for $E_{\nu}$ values close to $T_{p}$, where statistical errors increase. The most accurate predictions are obtained around $T_{p}\sim 100~\mathrm{GeV}$, corresponding to the energy range covered by NA49 and NA61 data for $\pi^{\pm}$ production. Both $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ receive contributions from $\pi^{+}$ and $\pi^{-}$ decays, as both pion charge can produce a $\nu_{\mu}$. In contrast, $\nu_{e}$ and $\bar{\nu}_{e}$ are primarily produced by $\pi^{+}$ and $\pi^{-}$ decays, respectively. As previously noted in ODDK22, the modeling of $\pi^{-}$ production carries slightly larger uncertainties due to less precise experimental data, leading to $\bar{\nu}_{e}$ having the largest model uncertainty among all $\nu$ flavors.
@@ IV Contribution from other production channels and from nuclei
In this section, we present our model for $\nu$ production from all the production channels other than the $\pi^{\pm}$ ones discussed in Sect. III. We also remind the tools for treating scatterings involving nuclei heavier than hydrogen.
@@ IV.1 Subdominant channels
Figure 4: Comparison among our differential cross sections and the one reported in Kamae et al. (2006) (Kamae), Bhatt et al. (2020) (Bhatt) and Koldobskiy et al. (2021) (AAfrag) as a function of $E_{\nu}$, for incident proton energy $T_{p}=10^{8},10^{6},10^{4}$ and $10^{2}$ GeV.
@@ IV.2 Production from heavier nuclei
For the inclusion of scatterings involving nuclei heavier than hydrogen, either in the CRs or in the ISM, we closely follow the prescriptions derived in ODDK22 for $\pi^{\pm}$. Specifically, when a $\pi^{\pm}$ is produced in collisions between a projectile and a target nucleus with mass numbers $A_{1}$ and $A_{2}$, the $\sigma_{\rm inv}$ formulas reported in Eqs.(7)–(10) of ODDK22 are modified according to the prescriptions in Eqs.(25)–(27) of ODDK22. The parameters in Eq.(26) are taken from the columns 1 and 2 Table V, where the columns for $\pi^{+}$ ($\pi^{-}$) are used to correct $\sigma_{\rm inv}$. The $K^{\pm}$ channel is modified analogously using columns 3 and 4 of Tab. V in ODDK22. For all other channels, we apply a correction function given by the average of the $K^{+}$ and $K^{-}$ cases.
@@ V Results on the $\nu$ production cross section and emissivity
Figure 5: Relative difference between our differential cross section and the one reported in Kamae et al. (2006) (Kamae), Bhatt et al. (2020) (Bhatt) and Koldobskiy et al. (2021) (AAfrag), for incident proton energy $T_{p}=10^{2},10^{4},10^{6}$ and $10^{8}$ GeV for $\nu_{\mu}$. We also show the statistical uncertainty of our model as a gray band.
We have now all the tools to compute the total differential cross section $d\sigma/dE_{\nu}$ for the inclusive production of neutrinos in inelastic $p+p$ collisions. The result is obtained by summing all contributions from $\pi^{\pm}$ and the subdominant channels, as discussed in Sections III and IV. These cross sections constitute the main result of our paper and are shown in Fig. 3 for a representative proton energy of $T_{p}=10^{6}$ GeV, for all flavors of neutrinos and antineutrinos. The $\pi^{\pm}$ channels dominate at all $\nu$ energies and for all flavors. All kaons contribute about 20% of the pions and 15% of the total, while all other subdominant channels scrape the 1-2% of the total. We have verified that these patterns are maintained across different proton energies. The gray curve and shaded band show the total $d\sigma/dE_{\nu}$ and its $1\sigma$ uncertainty, respectively. The resulting uncertainty spans from $6\%$ to $20\%$ depending on $E_{\nu}$, and is primarily driven by the modeling of the $\pi^{\pm}$ cross section, with larger uncertainties for $\bar{\nu}_{e}$ due to the dominant production of $\pi^{-}$, as explained in Section III. We remind that our cross-section model is validated at high $\sqrt{s}$ by a fit to ALICE and CMS data probing the midrapidity region only. This ensures that we capture the correct dependence on $\sqrt{s}$ of the pion multiplicity, which is predominantly produced at midrapidity. However, the spectral shape of our cross section—depending on $p_{T}$ and $x_{R}$—relies on extrapolation at high $\sqrt{s}$.
@@
Figure 6: The $\nu$ emissivity is computed for $p+p$, He+p+p, p+p+He, He+He and CNO+p+p scatterings for the different neutrino and antineutrino flavors. The gray line is the sum of all contributions (see text for details). Each prediction is plotted with the relevant uncertainty due to the production cross sections derived in this paper. In each bottom panel, the relative uncertainty to the total $\epsilon(E_{\nu})$ is displayed. For comparison, we show the results by Kamae et al. (2006) (Kamae), Bhatt et al. (2020) (Bhatt) and Koldobskiy et al. (2021) (AAfrag) in the $p+p$ channel.
In contrast, the Kamae et al. model shows deviations up to 50%, particularly at low energies, and displays significantly different behaviours across the flavors, with the exception of $\nu_{e}$, which remains close to the other models. We do not report Kamae model for $T_{p}=10^{6},10^{8}$ GeV, since they do not provide the tables for these energies. This discrepancy could be particularly relevant for the observation of a significant Galactic diffuse neutrino emission made by the IceCube Neutrino Observatory in Abbasi et al. (2023), where their template model is based on Kamae’s cross sections.
@@
The statistical uncertainty on $\epsilon(E_{\nu})$ due to hadronic production cross sections is approximately $5\%$ for $E_{\nu}\lesssim 10^{3}$ GeV, and increases to $10\%$ for $E_{\nu}\gtrsim 10^{6}$ GeV. For comparison, we show the results in dashed lines from Kamae et al. (2006) (Kamae), Koldobskiy et al. (2021) (AAfrag), and Bhatt et al. (2020) (Bhatt) for the $p+p$ channel. The models by Kamae and Bhatt do not provide full coverage across the proton and neutrino energy range required for this plot, and are therefore cut accordingly. There is excellent agreement at low energies ($E_{\nu}\lesssim 100$ GeV) among our model, AAfrag, and Bhatt. Overall, our results and Bhatt’s (orange line) agree within $\sim 10\%$, while AAfrag shows deviations up to $\sim 20\%$. Conversely, Kamae exhibits significant discrepancies below $10^{3}$ GeV (except for $\nu_{e}$), while remaining broadly consistent with the other models between $10^{3}$ GeV and $10^{4}$ GeV. As a general trend, our new cross sections predict an emissivity that, at least for the dominant $p+p$ channel, is a bit higher that the Monte Carlo state of the art.

### v2

@@ Abstract
The flux of astrophysical neutrinos is now measured with unprecedented accuracy and over several decades of energy spectrum. Their origin traces back to hadronic collisions between protons and nuclei in the cosmic rays with hydrogen and helium in the target gas. To accurately interpret the data, a precise determination of the underlying cross sections is therefore mandatory. We present a new evaluation of the neutrino production cross section from $p+p$ collisions, building on our previous analysis of the production cross section for $\pi^{\pm}$, $K^{\pm}$, and minor baryonic and mesonic channels. Cross sections for scatterings involving nuclei heavier than protons are also derived. The novelty of our approach is the analytical description of the Lorentz invariant cross section $\sigma_{\rm inv}$ and the fit of the model to the available accelerator data. We work with neutrino energies from $10$ GeV to $10^{7}$ GeV and, correspondingly, to incident proton (nuclei) energies from $10$ GeV to $10^{9}$ GeV (GeV/n). We obtain the total differential cross section $d\sigma(p+p\rightarrow\nu+X)/dE_{\nu}$ as a function of neutrino and proton energies, with an estimated uncertainty of 5% for neutrino energies below 100 GeV, increasing to 10% above TeV energies. Predictions are given for $\nu_{e},\nu_{\mu},\bar{\nu_{e}}$ and $\bar{\nu_{\mu}}$. A comparison with state-of-the-art cross sections, all relying on Monte Carlo generators, is also presented. To facilitate the use by the community, we provide numerical tables and a script for accessing our energy-differential cross sections.
@@ I Introduction
In this paper, we present a new accurate model that relies primarily on an analytic prescription fitted to the available experimental data, with the aim of determining their correct dependence on kinematic variables and of robustly quantifying the modeling uncertainties. The dominant production channel for $\nu_{e}$, $\nu_{\mu}$, $\bar{\nu_{e}}$ and $\bar{\nu_{\mu}}$ is the decay of $\pi^{\pm}$ mesons, for which we build upon the analysis performed in ODDK22. In addition, we carefully model the production cross sections of $K$ mesons and $\Lambda$ baryons, which contribute as well to the $\nu$ flux through their decay into $\pi^{\pm}$. Our strategy closely follows the approach adopted in ODDK22 for deriving cross sections relevant to the secondary production of CR electrons and positrons. While the parameterization of the individual channel cross sections were derived in ODDK22  the main innovation of this paper lies in the integration of these channels into a complete, data-driven model for the inclusive neutrino production cross section, along with a systematic quantification of the associated uncertainties.
@@ III $\nu$ from $p+p\rightarrow\pi^{\pm}+X$ collisions
The $\pi^{\pm}$ channel for $\nu$ production corresponds to the first two rows in Fig. 1(left) for $\nu_{\mu}$ and to the first row in Fig. 1(right) for $\nu_{e}$. Given its relevance, it is crucial to obtain precise data covering a wide kinematic phase space for the reaction $p+p\rightarrow\pi^{\pm}+X$. As discussed in ODDK22, measurements of $\sigma_{\rm inv}(p+p\rightarrow\pi^{\pm}+X)$ have been collected by various accelerator and collider experiments, spanning large portions of the kinematic phase space. In particular, NA49 Alt et al. (2005) and NA61 Aduszkiewicz et al. (2017) provide data at low $\sqrt{s}$, while ALICE Aamodt (2011) and CMS Chatrchyan et al. (2012); Sirunyan et al. (2017) cover the high-$\sqrt{s}$ regime. In this work we do not perform a new fit to the particle production data (since no new data have appeared); instead, we use the results for $\sigma_{\rm inv}$ obtained in ODDK22 to start the calculation of $\nu$ yields. The expressions for $\sigma_{\rm inv}$ are given in Eqs. (7)–(10) and the parameters for $\pi^{\pm}$ are listed in Table II of ODDK22 and in the Appendix A. We remark that the fit performed in ODDK22 is based on an analytical formula for $\sigma_{\rm inv}$, directly calibrated on the data and their uncertainties. The highest projectile energy on which we tested our model is $T_{p}=9\times 10^{7}$ GeV, corresponding to $\sqrt{s}=13$ TeV pion data from CMS Sirunyan et al. (2017). Beyond this limit, our model can only be extrapolated. However, care must be taken when treating purely statistical uncertainties at high $\sqrt{s}$. As explained in ODDK22, our model of $\sigma_{\rm inv}$ relies on a fit on NA49 and NA61 $\pi^{\pm}$ data at $\sqrt{s}=17.3$ GeV, which provide broad coverage in both $p_{T}$ and $x_{R}$, thereby constraining in a solid way the dependence of the model on these variables. The scaling of the cross section at higher energies was instead determined by fits to CMS and ALICE data at $\sqrt{s}>900$ GeV (corresponding to $T_{p}=4.3\times 10^{5}$ GeV in the LAB frame). These measurements, however, were performed at mid-rapidity and cover only a limited range in $p_{T}$, with no data available at forward rapidities. Consequently, our model implies an extrapolation of the cross section in the forward-rapidity region at high $\sqrt{s}$. To quantify the impact of this incomplete phase-space coverage, we will compare our results with predictions from other state-of-the-art Monte Carlo models, thereby estimating the overall uncertainties across different cross-section parametrizations.
@@
In Fig. 2, we show the differential cross section for the production of $\nu_{\mu}$, $\bar{\nu}_{\mu}$, $\nu_{e}$, and $\bar{\nu}_{e}$ from $\pi^{\pm}$ decays in $p+p$ collisions. Results are provided for different incident kinetic proton energies $T_{p}$ as a function of $E_{\nu}$, for all $\nu$ flavors. The uncertainties range from $5\%$ to $20\%$ over most of the energy range, except for $E_{\nu}$ values close to $T_{p}$, where statistical errors increase. The most accurate predictions are obtained around $T_{p}\sim 100~\mathrm{GeV}$, corresponding to the energy range covered by NA49 and NA61 data for $\pi^{\pm}$ production. Both $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ receive contributions from $\pi^{+}$ and $\pi^{-}$ decays, as both pion charges can produce a $\nu_{\mu}$. In contrast, $\nu_{e}$ and $\bar{\nu}_{e}$ are primarily produced by $\pi^{+}$ and $\pi^{-}$ decays, respectively. As previously noted in ODDK22, the modeling of $\pi^{-}$ production carries slightly larger uncertainties due to less precise experimental data, leading to $\bar{\nu}_{e}$ having the largest model uncertainty among all $\nu$ flavors.
@@ IV Contribution from other production channels and from nuclei
In this section, we present our model for $\nu$ production from all the production channels other than the $\pi^{\pm}$ ones discussed in Sect. III. We also remind the reader of the tools for treating scatterings involving nuclei heavier than hydrogen.
@@ IV.1 Subdominant channels
Figure 4: Comparison among our differential cross sections and the one reported in Kamae et al. (2006) (Kamae), Bhatt et al. (2020) (Bhatt), Koldobskiy et al. (2021) (AAfrag) and Kelner et al. (2006) (Kelner) as a function of $E_{\nu}$, for incident proton energy $T_{p}=10^{8},10^{6},10^{4}$ and $10^{2}$ GeV.
@@ IV.2 Production from heavier nuclei
For the inclusion of scatterings involving nuclei heavier than hydrogen, either in the CRs or in the ISM, we closely follow the prescriptions derived in ODDK22 for $\pi^{\pm}$. Specifically, when a $\pi^{\pm}$ is produced in collisions between a projectile and a target nucleus with mass numbers $A_{1}$ and $A_{2}$, the $\sigma_{\rm inv}$ formulas reported in Eqs. (7)–(10) of ODDK22 are modified according to the prescriptions in Eqs. (25)–(27) of ODDK22. The parameters in Eq. (26) are taken from the columns 1 and 2 in Table V of ODDK22, where the columns for $\pi^{+}$ ($\pi^{-}$) are used to correct $\sigma_{\rm inv}$. The $K^{\pm}$ channel is modified analogously using columns 3 and 4 of Tab. V in ODDK22. For all other channels, we apply a correction function given by the average of the $K^{+}$ and $K^{-}$ cases.
@@ V Results on the $\nu$ production cross section and emissivity
Figure 5: Relative difference between our differential cross section and the one reported in Kamae et al. (2006) (Kamae), Bhatt et al. (2020) (Bhatt), Koldobskiy et al. (2021) (AAfrag) and Kelner et al. (2006) (Kelner), for incident proton energy $T_{p}=10^{2},10^{4},10^{6}$ and $10^{8}$ GeV for $\nu_{\mu}$. We also show the statistical uncertainty of our model as a gray band.
We have now all the tools to compute the total differential cross section $d\sigma/dE_{\nu}$ for the inclusive production of neutrinos in inelastic $p+p$ collisions. The result is obtained by summing all contributions from $\pi^{\pm}$ and the subdominant channels, as discussed in Secs III and IV. These cross sections constitute the main result of our paper and are shown in Fig. 3 for a representative proton energy of $T_{p}=10^{6}$ GeV, for all flavors of neutrinos and antineutrinos. The $\pi^{\pm}$ channels dominate at all $\nu$ energies and for all flavors. All kaons contribute about 20% of the pions and 15% of the total, while all other subdominant channels scrape the 1-2% of the total. We have verified that these patterns are maintained across different proton energies. The gray curve and shaded band show the total $d\sigma/dE_{\nu}$ and its $1\sigma$ uncertainty, respectively. The resulting uncertainty spans from $6\%$ to $20\%$ depending on $E_{\nu}$, and is primarily driven by the modeling of the $\pi^{\pm}$ cross section, with larger uncertainties for $\bar{\nu}_{e}$ due to the dominant production of $\pi^{-}$, as explained in Sec III. We remind that our cross-section model is validated at high $\sqrt{s}$ by a fit to ALICE and CMS data probing the midrapidity region only. This ensures that we capture the correct dependence on $\sqrt{s}$ of the pion multiplicity, which is predominantly produced at midrapidity. However, the spectral shape of our cross section—depending on $p_{T}$ and $x_{R}$—relies on extrapolation at high $\sqrt{s}$.
@@
Figure 6: The $\nu$ emissivity is computed for $p+p$, He+p+p, p+p+He, He+He and CNO+p+p scatterings for the different neutrino and antineutrino flavors. The gray line is the sum of all contributions (see text for details). Each prediction is plotted with the relevant uncertainty due to the production cross sections derived in this paper. In each bottom panel, the relative uncertainty to the total $\epsilon(E_{\nu})$ is displayed. For comparison, we show the results by Kamae et al. (2006) (Kamae), Bhatt et al. (2020) (Bhatt), Koldobskiy et al. (2021) (AAfrag) and Kelner et al. (2006) (Kelner) in the $p+p$ channel.
In contrast, the Kamae et al. model shows deviations up to 50%, particularly at low energies, and displays significantly different behaviours across the flavors, with the exception of $\nu_{e}$, which remains close to the other models. We do not report Kamae model for $T_{p}=10^{6},10^{8}$ GeV, since they do not provide the tables for these energies. This discrepancy could be particularly relevant for the observation of a significant Galactic diffuse neutrino emission made by the IceCube Neutrino Observatory in Abbasi et al. (2023), where their template model is based on Kamae’s cross sections. In the end, the parametrization by Kelner Kelner et al. (2006)—which does not provide separate cross sections for each neutrino and antineutrino species but instead gives averaged cross sections for $\nu_{e}$$\bar{\nu}_{e}$ and $\nu_{\mu}$$\bar{\nu}_{\mu}$—shows significant deviations from all other models, especially at high $T_{p}$. It predicts a noticeably harder dependence on $E_{\nu}$ compared to the other approaches, with particularly large discrepancies for $\nu_{e}$ and $\bar{\nu}_{e}$. In this case, using an average between the two species introduces substantial errors, especially at low energies where the $\pi^{\pm}$ production cross sections differ significantly.
@@
The statistical uncertainty on $\epsilon(E_{\nu})$ due to hadronic production cross sections is approximately $5\%$ for $E_{\nu}\lesssim 10^{3}$ GeV, and increases to $10\%$ for $E_{\nu}\gtrsim 10^{6}$ GeV. For comparison, we show the results in dashed lines from Kamae et al. (2006) (Kamae), Koldobskiy et al. (2021) (AAfrag), and Bhatt et al. (2020) (Bhatt) and Kelner et al. (2006) (Kelner) for the $p+p$ channel. The models by Kamae and Bhatt do not provide full coverage across the proton and neutrino energy range required for this plot, and are therefore cut accordingly. There is excellent agreement at low energies ($E_{\nu}\lesssim 100$ GeV) among our model, AAfrag, and Bhatt. Overall, our results and Bhatt’s (orange line) agree within $\sim 10\%$, while AAfrag shows deviations up to $\sim 20\%$. Conversely, Kamae exhibits significant discrepancies below $10^{3}$ GeV (except for $\nu_{e}$), while remaining broadly consistent with the other models between $10^{3}$ GeV and $10^{4}$ GeV. Kelner’s model shows significant discrepancies (e.g. up to 50% or larger for $E_{\nu}>20$ GeV) at all energies and across all channels. The particularly large mismatch for the $\bar{\nu}_{e}$ channel (up to 100% for $E_{\bar{\nu}_{e}}>20$ GeV) arises because the model uses an average of the $\pi^{-}$ and $\pi^{+}$ cross sections, which strongly overestimates the $\bar{\nu}_{e}$ yield. Overall, the Kelner parametrization substantially overestimates neutrino production in every channel, as already noted in Koldobskiy et al. (2021). As a general trend, our new cross sections predict an emissivity that, at least for the dominant $p+p$ channel, is a bit higher that the Monte Carlo state of the art.